# Radiant Triangle Soup with Soft Connectivity Forces for 3D Reconstruction and Novel View Synthesis

### v1

@@ Abstract
In this work, we introduce an inference-time optimization framework utilizing triangles to represent the geometry and appearance of the scene. More specifically, we develop a scene optimization algorithm for triangle soup, a collection of disconnected semi-transparent triangle primitives. Compared to the current most-widely used primitives for 3D scene representation, namely Gaussian splats, triangles allow for more expressive color interpolation, and benefit from a large algorithmic infrastructure for downstream tasks. Triangles, unlike full-rank Gaussian kernels, naturally combine to form surfaces. We formulate connectivity forces between triangles during optimization, encouraging explicit, but soft, surface continuity in 3D. We perform experiments on a representative 3D reconstruction dataset and show competitive photometric and geometric results.
@@ 1 Introduction
Gaussian splatting has shown impressive results in Novel View Synthesis (NVS) and 3D reconstruction. However 3DGS [1] and the subsequent work fail to accurately model the geometry of surfaces. 3D Gaussians are by design diffuse volumetric primitives, which inherently causes difficulties in representing geometry and sharp boundaries. This is because they are optimized for novel view synthesis, which is achieved by alpha-blending several overlapping Gaussians, none of which has to be located on the underlying surfaces. In order to compensate, optimizing over these volumetric primitives leads to an expensive increase in the number of primitives and can cause floaters or blurry artifacts in the scene.
To rectify this issue, Gaussian Surfels [2, 3, 4] and General 3D Convexes [5] have been proposed to help with accuracy in geometry reconstruction. Gaussian Surfels use either strict 2D Gaussians [2] or flatten 3D Gaussians via loss regularization [3, 4] to better model thin surfaces, but these diffuse ellipsoids can still fail to accurately represent surfaces without substantially increasing their density. For clarity, throughout this paper we will refer to the work of Kerbl et al. [1] and Huang et al. [2] as 3DGS and 2DGS, respectively. We will use Volumetric Gaussian Splatting (VGS) and Surfel Gaussian Splatting (SGS) in general without referring to a particular approach.
@@
In this paper, we present a new, complete methodology for 3D reconstruction that uses semi-transparent triangle primitives as the scene representation for Radiance Fields (RF). Our approach, which we named Radiant Triangle Soup (RTS), includes non-differentiable mechanisms for initialization, pruning, and densification, as well as differentiable mechanisms for rasterization. Our loss comprises 2D and 3D terms that favor more accurate rendering of color and surface normals as well as smoother surfaces. Within this framework, we have implemented an explicit Radiance Field algorithm and have achieved competitive results in terms of both appearance and geometry (See Fig. 1).
@@
RTS provides a feature that is not supported by current GS scene representations, in which there is no avenue for explicit information sharing in 3D among primitives. Instead, Gaussian kernels communicate via alpha-blending when rendered onto common pixels. Back-propagating from the loss at each pixel to the primitives is the only means of coordination across Gaussians. Using multiple images from various viewpoints gives rise to several indirect constraints on the primitives, but direct constraints are currently absent. Our approach, on the other hand, is the first to enable direct information exchange among neighboring primitives. Having explicit 3D communication allows for a more direct coordination of primitive behavior throughout the optimization process.
@@
Furthermore, most of the scene representations thus far operate with a single color parametrization for each primitive. This can lead to large numbers of overlapping primitives in order to reconstruct details in the surfaces of the scene. Using triangles as a representation allows for each triangle vertex to encode a separate color, leading to more expressive primitives through bilinear interpolation.
@@
  • Develops a new SGS scene representation through alpha-blending of triangle primitives.
@@
  • Incorporates a depth smoothness loss from the unsupervised Multi-View Stereo (MVS) literature.
Refer to caption Refer to caption Refer to caption
Figure 1: Optimizing over Radiant Triangle Soup (RTS) produces high quality 3D models from captured images. Above are the direct visualization of the point clouds without meshing from the DTU dataset [6]. See Section 6 for details.
@@ 2 Related Work
We begin this section with Gaussian splatting formulations that favor surfaces and continue with methods that rely on non-Gaussian primitives. Surveys on other aspects of Gaussian splatting, that we cannot cover due to space limitations, include [7, 8, 9, 10]
Gaussian Splatting
@@
The seminal work of Kerbl et al. on 3D Gaussian Splatting [1] introduced an explicit alternative to NeRF [11] that is able to achieve high-quality rendering at much higher speed. 3DGS relies on interleaved differentiable optimization and non-differentiable density control of the explicit representation. The objective function of the former guides the optimization process to decrease view synthesis errors for one of the training images at each iteration.
@@
2D Gaussian Splatting (2DGS) [2] modified 3DGS to prioritize the reconstruction of surfaces, rather than volumetric material. This was accomplished by collapsing the 3D GS to 2D discs by setting the minimum eigenvalue of the Gaussian to 0. Similar approaches were developed by the authors of SuGaR [3] and PGSR [4] who devised techniques for aligning 2D Gaussian with the surfaces via regularization and a multi-view loss, respectively. Gaussian Surfels [12] are also similarly relying on monocular surface normal estimates and a normal-depth consistency loss. Gaussian Opacity Fields (GOF) [13] enable direct geometry extraction as the zero-level set of 3D Gaussians without the need for external surface extraction modules.
3D-Half-Gaussian Splatting (3D-HGS) [14] enables the representation of perfectly planar surfaces and hard edges by attaching a splitting plane to each Gaussian and setting the density of one of the half-spaces to 0. The high-density splitting plane can then be aligned with the local surface.
Alternative Representations
Representations based on NeRF and 3DGS have found great success, but have also inspired researchers to seek alternatives. NeuRBF [15] utilizes Radial Basis Functions (RBFs) to overcome limitations of NeRF due to the global nature of its MLP and features. Instant-NGP [16] and Plenoxels [17] proposed representations with local features, but they were arranged in regular grids. In contrast, NeuRBF achieves spatial adaptivity and more precise handling of high frequency content via the use of anisotropic RBFs.
Recently, alternative primitives have begun to emerge. DARB-Splatting [18] is motivated by the observation that a broader class of decaying anisotropic RBFs (DARBFs), than just Gaussians, have 3D-to-2D projections that can be integrated easily. Compared to Gaussians, other DARBFs lead to faster convergence with a smaller memory footprint. GES [19] replaces the Gaussian kernel with a Generalized Exponential Function, overcoming the low-pass effect of the Gaussian and thus requiring far fewer primitives to represent the scene. Motivated by artifacts caused by the opacity of Gaussian splats falling off away from the centroid, SolidGS [20] modifies the Gaussian kernel by introducing a learnable solidness factor that makes the splats exhibit a more solid-like behavior. Chen et al. [21] argue that unlike Gaussian kernels that have unbounded support and may cause blurring, bounded linear kernels reduce foreground-background interference and yield clearer and more accurate reconstructions.
@@
Besides the methods that force their Gaussians to be planar [2, 3, 4], but still diffuse, there are others that represent the scene with collections of planar primitives. Zanjani et al. [22] initially use Gaussian splats to model the scene and then merge them hierarchically to infer 3D planes, which are abundant in indoor scenes. PlanarSplatting [23] also addresses surface reconstruction in indoor scenes. In this case, planes are the only primitives, initialized via monocular depth estimation. TRIPS [24] uses points as primitives. The key principle is that points can be rasterized into an image pyramid from which the appropriate layer can be selected according to the size of the projected point. Holes can be filled by a small network yielding accurate, crisp renderings.
@@
A few methods rely on volumetric primitives. LinPrim [25] is based on linear solid primitives, specifically octahedra and tetrahedra, with triangular faces that take advantage of well known fast processing and rendering mechanisms. LinPrim performs volumetric rendering by computing the entrance and exit points of each ray through the primitives. 3D Convex Splatting (3DCS) [5] was inspired by the same shortcoming of GS: very large numbers of primitives are needed to model hard edges and flat surfaces due to the diffuseness of the Gaussians. As the name suggests, the scene is represented by a set of polyhedral convexes which undergo volumetric rendering, pruning and splitting operations. The concepts of smoothness and sharpness introduced by 3DCS have inspired our diffuseness (see Section 4). Radiant Foam [26] is both efficient, like GS, but also enables modeling light transport phenomena like reflection and refraction. The method takes its name from the representation which is a dense Voronoi tessellation of 3D space. The positions of Voronoi sites are differentiable after rendering, allowing iterative optimization. Since this is the only representation with connected elements, non-differentiable operations, e.g. pruning, have to be implemented carefully. All methods in this paragraph rely on different forms of piecewise constant volumetric elements with planar faces. Computing the intersections of each primitive with a ray requires multiple ray-triangle intersections, and is thus more expensive than our method.
@@
A few common themes emerge by analyzing the above methods. Most advocate the use of bounded primitives to enhance their ability to represent sharp edges and flat surfaces with small numbers of primitives. Our primitives are bounded but have diffuse boundaries to facilitate optimization. Like SolidGS [20], the opacity of our triangles remains approximately constant inside the triangle leading to sharper synthesized views. Several of the non-Gaussian primitives [5, 23, 25] leverage the existing software and hardware infrastructure by proposing representations with triangular faces.
3 Background
Before discussing our optimization framework, we first introduce the necessary background geometry used to parameterize a ray-triangle intersection in barycentric coordinates. The barycentric coordinates of a point on a triangle $\lambda=[\lambda^{0},\lambda^{1},\lambda^{2}]^{T}$ are one form of expressing the coordinates of the point as a linear combination of the coordinates of the three vertices, $[V^{0},V^{1},V^{2}]$. This parameterization is important for graphics applications to be able to efficiently rasterize triangles onto the screen space and interpolate the color from each vertex for a pixel. Throughout this paper, we follow the assumption that $\lambda^{j}\in\mathbb{R}^{+}$ corresponds to vertex $V^{j}\in\mathbb{R}^{3}$ such that $\sum_{j=1}^{3}\lambda^{j}=1$.
Given a pixel, we compute the ray-triangle intersection,
$P=C+(\widehat{r}d)$ (1)
$d=\frac{\widehat{n}\cdot\overrightarrow{CB}}{\widehat{n}\cdot\widehat{r}}$ (2)
$\lambda=\frac{1}{\widehat{n}\cdot(\overrightarrow{V_{0}V_{1}}\times\overrightarrow{V_{0}V_{2}})}\begin{bmatrix}\widehat{n}\cdot(\overrightarrow{V_{1}V_{2}}\times\overrightarrow{V_{1}P})\\ \widehat{n}\cdot(\overrightarrow{V_{2}V_{0}}\times\overrightarrow{V_{2}P})\\ \widehat{n}\cdot(\overrightarrow{V_{0}V_{1}}\times\overrightarrow{V_{0}P})\end{bmatrix}$ (3)
@@ 4 Method
Given a set of images together with camera poses and a set of sparse points $\mathcal{S}$ computed via Structure-From-Motion [27], we construct an explicit scene representation using triangle primitives, the parameterization of which we discuss in Section 4.1. To model the surfaces in the scene, we render into each camera the image, depth, and normal map of the triangles through alpha-blending (see Section 4.4). The triangle parameters are directly updated via back-propagation after computing losses between the rendering of the scene and the ground truth image from each view. We also include 2D losses on the rendered depth and normal maps (see Section 5). In our representation, primitives maintain soft connectivity with their neighbors, discussed in Section 4.5, for which we compute additional loss directly over the connection orientations (see Section 5). Similar to previous works [2, 3, 4, 1, 13], we develop an adaptive densification strategy in order to facilitate the addition and removal of primitives in the scene (see Section 4.6). We show an overview of our optimization algorithm in Fig. 2.
@@ 4.1 Parameterization
where $\mu$ is the triangle reference point, $\Delta=\begin{bmatrix}\delta^{0}&\delta^{1}&\delta^{2}\end{bmatrix}$ are the per-vertex Spherical Harmonics, $s=\begin{bmatrix}s^{0}&s^{1}&s^{2}\end{bmatrix}$ are the vertex scales, $R$ is the $3\times 3$ rotation matrix, $\alpha$ is the opacity, and $\sigma$ is the diffuse scalar, discussed in Section 4.3.
@@ 4.2 Initialization
The coordinates of the sparse SfM points are used as the initialization for the triangle reference points, and the point colors as the initialization for the zero-component of the Spherical Harmonics for all three vertices. Similar to previous works, the scale for our primitives is computed based on the average distance to the three nearest neighboring points. Each triangle starts as an equilateral triangle, using the scale to parameterize the distance of each vertex from the triangle reference point $\mu$ (see Fig. 3). Note that in this representation, the aforementioned triangle reference point $\mu$ is not the barycenter of the triangle; that is, $\mu$ is not guaranteed to be the mean of all three vertices, but instead represents the anchoring position of the triangle. The diffuse scalar $\sigma$ is also set as a function of the average distance to the three nearest neighboring points. The parameterization of the initial rotation matrix for each triangle is defined such that the $\overrightarrow{V^{2}V^{0}}$ edge and the triangle normal $\widehat{n}$ are used as the $x-$ and $z-$ axis, respectively. We then initialize each triangle by applying a random rotation.
Figure 3: Triangle parameterization. Each triangle is parameterized by a reference point $\mu$, and three scale values $\begin{bmatrix}s^{0}&s^{1}&s^{2}\end{bmatrix}$. These parameters, along with the rotation matrix $R$, define the coordinates of each vertex $V^{j}$.

### v2

@@ Abstract
We introduce an inference-time scene optimization algorithm utilizing triangle soup, a collection of disconnected translucent triangle primitives, as the representation for the geometry and appearance of a scene. Unlike full-rank Gaussian kernels, triangles are a natural, locally-flat proxy for surfaces that can be connected to achieve highly complex geometry. When coupled with per-vertex Spherical Harmonics (SH), triangles provide a rich visual representation without incurring an expensive increase in primitives. We leverage our new representation to incorporate optimization objectives and enforce spatial regularization directly on the underlying primitives. The main differentiator of our approach is the definition and enforcement of soft connectivity forces between triangles during optimization, encouraging explicit, but soft, surface continuity in 3D. Experiments on representative 3D reconstruction and novel view synthesis datasets show improvements in geometric accuracy compared to current state-of-the-art algorithms without sacrificing visual fidelity.
Figure 1: Optimizing Radiant Triangle Soup (RTS) produces high quality 3D models from captured images. Above, from left to right, are the rendered normal map, rendered depth map, rendered image, and the direct rasterization of the triangle primitives, as if they were opaque.
@@ 1 Introduction
Gaussian Splatting (GS) methods are effective at Novel View Synthesis (NVS). However 3D Gaussian Splatting (3DGS) (Kerbl et al., 2023) fails to accurately model the geometry of scene surfaces. 3D Gaussians are by design smooth, unbounded volumetric primitives, which are inherently ill-suited for representing flat surfaces and sharp boundaries. Optimization for novel view synthesis involves alpha-blending several overlapping Gaussians, none of which have to be located on the underlying surfaces. As a result, optimizing a representation of 3D Gaussian kernels for image synthesis commonly causes floaters or blurry artifacts in the scene.
Several authors (Huang et al., 2024; Guédon & Lepetit, 2024; Chen et al., 2024a; Dai et al., 2024) have proposed flattening the kernels, with either strict 2D Gaussians (Huang et al., 2024) or flattened 3D Gaussians via loss regularization (Guédon & Lepetit, 2024; Chen et al., 2024a), to better model thin surfaces. These methods have shown that modifying the underlying primitives and enforcing optimization objectives on rendered geometry leads to improvements in geometric accuracy. While flat primitive help with surface alignment, diffuse ellipsoids still fail to accurately model depth discontinuities (see Fig. 8).
Triangles are the fundamental primitives in computer graphics because they are necessarily planar and convex. They can approximate any surface as piecewise planar at any tolerance level, can directly model sharp discontinuities, and can be rendered rapidly without approximations.
@@
We present a new methodology for 3D reconstruction, which we named Radiant Triangle Soup (RTS) that enables gradient-based optimization of a Radiance Field (RF) using translucent triangle primitives as the scene representation (see Fig. 1). We provide a complete framework, including differentiable mechanisms for rasterization, as well as non-differentiable mechanisms for initialization, pruning, and densification. Our experiments show RTS achieves competitive results in terms of both appearance and geometry (see Fig. 7).
@@
Conceptually, RTS provides a feature that is not supported by any other GS scene representation: an avenue for explicit information sharing among 3D among primitives. Conventional Gaussian kernels interact via alpha-blending when rendered onto common pixels. Back-propagating from the loss at each pixel to the primitives is the only means of coordination across Gaussians. Using multiple images from various viewpoints gives rise to several indirect constraints on the primitives, but direct constraints are currently absent.
Conversely, RTS is the first framework to enable direct information exchange among neighboring primitives. We formulate connectivity losses between neighboring triangle edges, allowing for a more direct and effective coordination of, and constraint on, primitive behavior throughout the optimization process. Even though using triangles as primitives alone achieves better geometric accuracy compared to Gaussian surfels, encouraging connectivity during optimization leads to more accurate surfaces with less floating artifacts (see Table 3).
@@
Color expressivity on a per-primitive basis is another dimension where the RTS representation excels. It is common practice to use a single color parametrization for each primitive. However, this may require large numbers of overlapping primitives in order to reconstruct details in the surfaces of the scene. Using triangles as the representation allows for each vertex to encode a separate color (in the form of Spherical Harmonics), leading to more expressive primitives through bilinear interpolation.
@@
  • Develops a new scene representation through alpha-blending of triangle primitives.
@@
  • Increases primitive expressivity via multi-color encoding.
@@ 2 Related Work
We begin this section with Gaussian splatting formulations that favor surfaces and continue with methods that rely on non-Gaussian primitives. Surveys on other aspects of Gaussian splatting, omitted due to space limitations, include (Bao et al., 2025; Dalal et al., 2024; Luo et al., 2024; Wu et al., 2024). The seminal work of Kerbl et al. (2023) on 3D Gaussian Splatting introduced an explicit alternative to NeRF (Mildenhall et al., 2020) that is able to achieve high-quality rendering at much higher speed. 3DGS relies on interleaved differentiable optimization and non-differentiable density control of the explicit representation. The optimization process decreases view synthesis errors for one of the training images at each iteration and density control guides the placement of primitives.
@@
2D Gaussian Splatting (2DGS) (Huang et al., 2024) modified 3DGS to prioritize the reconstruction of surfaces, rather than volumetric material. This was accomplished by collapsing the 3D GS to 2D discs by setting the minimum eigenvalue of the Gaussian to 0. The authors of SuGaR (Guédon & Lepetit, 2024) and PGSR (Chen et al., 2024a) devised similar techniques for aligning 2D Gaussian with the surfaces via regularization and a multi-view loss, respectively. Gaussian Surfels (Dai et al., 2024) rely on monocular surface normal estimates and a normal-depth consistency loss. Gaussian Opacity Fields (GOF) (Yu et al., 2024) extract surfaces as the zero-level set of 3D Gaussians. 3D-Half-Gaussian Splatting (3D-HGS) (Li et al., 2024) enables the representation of perfectly planar surfaces and hard edges by attaching a splitting plane to each Gaussian, aligning it to the local surface, and setting the density of one of the half-spaces to 0.
Representations based on NeRF and 3DGS have found great success, but have also inspired researchers to seek alternatives. NeuRBF (Chen et al., 2023) utilizes Radial Basis Functions (RBFs) to overcome limitations of NeRF due to the global nature of its MLP and features. GES (Hamdi et al., 2024) is based on an explicit representation which replaces the Gaussian kernel with a Generalized Exponential Function, overcoming the low-pass effect of the Gaussian and thus requiring fewer primitives to represent the scene. Similar approaches based on smooth, non-Gaussian kernels include DARB-Splatting (Arunan et al., 2025), SolidGS (Shen et al., 2024), Beyond Gaussians (Chen et al., 2024b) and Deformable Beta Splatting (Liu et al., 2025).
@@
Besides the methods that force their Gaussians to be planar (Huang et al., 2024; Guédon & Lepetit, 2024; Chen et al., 2024a), but still diffuse, there are others that represent the scene with collections of planar primitives. Zanjani et al. (2025) initially use Gaussian splats to model the scene and then merge them into 3D planes, which are abundant in indoor scenes. PlanarSplatting (Tan et al., 2025) is also designed for indoor scenes using planes, initialized via monocular depth estimation, as the only primitives. TRIPS (Franke et al., 2024) is based on the principle that point primitives can be rasterized into an image pyramid from which the appropriate layer can be selected according to the size of the projected point. Holes can be filled by a small network yielding accurate, crisp renderings. Triangle Splatting (Held et al., 2025a) uses triangles with diffuse edges as primitives, but does not support any mechanism for them to interact directly with each other.
Non-planar primitives were introduced by BG-Triangle (Wu et al., 2025) which uses Bézier Gaussian triangles that are effective near boundaries, but comes at the cost of operating on complex, non-planar primitives that are hard to render. Quadratic Gaussian Splatting (Zhang et al., 2025) uses deformable quadratic surfaces as primitives and geodesic, instead of Euclidean, distance-based density distributions that adapt to the curvature of the primitives.
@@
Another class of methods rely on volumetric primitives. LinPrim (von Lützow & Nießner, 2025) is based on linear solid primitives with triangular faces and performs volumetric rendering by computing the entrance and exit points of each ray through the primitives. 3D Convex Splatting (3DCS) (Held et al., 2025b) was inspired by the limitation of GS that requires very large numbers of primitives to model hard edges and flat surfaces due to the diffuseness of the Gaussians. As the name suggests, the scene is represented by a set of polyhedral convexes which undergo volumetric rendering, pruning and splitting operations. The concepts of smoothness and sharpness introduced by 3DCS have inspired our diffuseness (see Section 3). Radiant Foam (Govindarajan et al., 2025) enables modeling light transport phenomena, like reflection and refraction, by tessellating the space into Voronoi cells and iteratively optimizing the positions of the Voronoi vertices. All methods in this paragraph rely on different forms of volumetric elements with planar faces. Computing the intersections of each primitive with a ray requires multiple ray-triangle intersections, and is thus more expensive than RTS.
A recent trend in the literature has been joint optimization of two representations: one for synthesizing novel views and one that is more faithful to the surfaces (Choi et al., 2024; Jiang et al., 2025). MILo (Guédon et al., 2025) maintains Gaussians that are alpha-blended for view synthesis and a watertight mesh without texture that captures the geometry of the scene.
@@
A few common themes emerge by analyzing the above methods. Most advocate the use of bounded primitives to enhance their ability to represent sharp edges and flat surfaces with small numbers of primitives. Our primitives are bounded but have diffuse boundaries to facilitate optimization. No other method, however, has a mechanism for direct inter-primitive communication like RTS.
@@ 3 Method
Figure 3: Triangle parameterization. Each triangle is parameterized by the incenter $\mu$, and three scale values $\begin{bmatrix}s^{0}&s^{1}&s^{2}\end{bmatrix}$. These parameters, along with the rotation matrix $R$, define the coordinates of each vertex $V^{j}$.
Given a set of images together with camera poses and a set of sparse points $\mathcal{S}$ computed via Structure-From-Motion (Schönberger & Frahm, 2016), we construct an explicit scene representation using triangle primitives, the parameterization and initiliazation of which we discuss in Sections 3.1 and 3.2, respectively. The triangles are endowed with diffuse boundaries, similar to 3DCS (Held et al., 2025b) (Section 3.3). To model the surfaces in the scene, we render into each camera the image, depth, and normal map of the triangles through alpha-blending (see Section 3.4). The triangle parameters are directly updated via back-propagation after computing losses between the rendering of the scene and the ground truth image from each view. We also include 2D losses on the rendered depth and normal maps (see Section 4). In our representation, primitives maintain soft connectivity with their neighbors, discussed in Section 3.5, for which we compute additional loss directly over the connection orientations (see Section 4). Similar to previous works (Kerbl et al., 2023; Huang et al., 2024; Guédon & Lepetit, 2024; Chen et al., 2024a; Yu et al., 2024), we develop a strategy for adaptive density control in order to facilitate the addition and removal of primitives in the scene (see Section 3.6). We show an overview of our optimization algorithm in Fig. 2.
@@ 3.1 Parameterization
where $\mu$ is the incenter of the triangle, $\Delta=\begin{bmatrix}\delta^{0}&\delta^{1}&\delta^{2}\end{bmatrix}$ are the per-vertex Spherical Harmonics, $s=\begin{bmatrix}s^{0}&s^{1}&s^{2}\end{bmatrix}$ are the vertex scales, $R$ is the $3\times 3$ rotation matrix, $\alpha$ is the opacity, and $\sigma$ is the diffuse scalar, discussed in Section 3.3.
@@ 3.2 Initialization
The coordinates of each sparse SfM point are used as the initialization for the incenter, and the point colors as the initialization for the zero-component of the Spherical Harmonics for all three vertices, which are then optimized separately. Similar to previous works, the scale for our primitives is computed based on the average distance to the three nearest neighboring points. Each triangle starts as equilateral, using the scale to parameterize the distance of each vertex from the incenter of the triangle $\mu$ along the bisectors of the angles (see Fig. 3). The diffuse scalar $\sigma$ is also set as a function of the average distance to the three nearest neighboring points. The parameterization of the initial rotation matrix for each triangle is defined such that the $\overrightarrow{V^{2}V^{0}}$ edge and the triangle normal $\widehat{n}$ are used as the $x$- and $z$- axis, respectively. We then initialize each triangle by applying a random rotation to this matrix.